### v1

@@ 7.2 Delayed Start
This section introduces a technique presented in [25]. We now fix a stochastic basis $\gamma$, an initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$, and suppress the dependence on $\gamma$ in this section. For $\alpha\in{\mathscr{A}}$ and $\epsilon>0$, we construct an approximating state process $X^{\alpha,\epsilon}$ and its cost. We first set
@@
$X^{\alpha,\epsilon}_{t}=X_{0}+\int_{\epsilon}^{t}$ $b(s,X^{\alpha,\epsilon}_{\cdot\wedge s},\alpha_{s})\,\mathrm{d}s+\int_{\epsilon}^{t}\boldsymbol{\sigma}(s,X^{\alpha,\epsilon}_{\cdot\wedge s},\alpha_{s})\,\mathrm{d}{\mathscr{W}}_{s}$
@@
Using ${\mathcal{L}}^{\alpha,\epsilon}_{t}:={\mathcal{L}}((X^{\alpha,\epsilon}_{\cdot\wedge t},\alpha_{t})\,|\,{\mathscr{G}})$, we define the approximating cost functional by,
@@
We first fix $\alpha\in{\mathscr{A}}$, $\epsilon>0$, and we suppress the dependence on $\mathbb{E}|X_{0}|^{2}$, and on the constants in Assumptions 2.3 and 7.1. To simplify the notation, we set
@@
$\mathbb{E}[\|X^{\alpha}-X^{\alpha,\epsilon}\|^{2}_{\epsilon}]=\mathbb{E}[\sup_{s\in[0,\epsilon]}|X^{\alpha}_{t}-X_{0}|^{2}]\leq c_{1}(\epsilon+\|\alpha\|^{2}_{2,\epsilon}).$
We then use the Burkholder-Davis-Gundy inequality, [42, Corollary 2.12], the Lipschitz property of the coefficients $(b,\boldsymbol{\sigma},\boldsymbol{\lambda})$, and the Grönwall’s inequality to arrive at
@@
$|F(t,X^{k}_{\cdot\land t},\alpha^{k}_{t},{\mathcal{L}}^{k}_{t})-F(t,Y^{k}_{\cdot\land t},\alpha^{k}_{t},\widehat{\mathcal{L}}^{k}_{t})|+|G(X^{k}_{\cdot\land T},{\mathcal{L}}^{k}_{T})-G(Y^{k}_{\cdot\land T},\widehat{\mathcal{L}}^{k}_{T})|\leq m(\zeta_{k}),$
for every $t\in[0,T]$, where $m$ is the modulus of continuity of $F$ and $G$. Consequently, the following sequence of random variables
@@ 7.3 Simple Controls
for some $t_{0}=0<t_{1}=:t_{1}(\alpha)<..<t_{m+1}=T$, ${\mathscr{F}}_{t_{i}}^{\gamma}$-measurable $\alpha_{i}:\Omega^{\gamma}\mapsto U$ and $a_{0}\in U$.
@@
${\mathcal{L}}^{\gamma}(X_{0},\alpha,\Xi^{\gamma,\epsilon}_{T})={\mathcal{L}}^{\tilde{\gamma}}(\tilde{X}_{0},\tilde{\alpha},\Xi^{\tilde{\gamma},\epsilon}_{T})\quad\Rightarrow\quad J_{g}^{\gamma,\epsilon}(X_{0},\alpha)=J_{g}^{\tilde{\gamma},\epsilon}(\tilde{X}_{0},\tilde{\alpha}).$
@@
Consider the measurable space ${\mathscr{Y}}:=\mathbb{R}^{n}\times U^{m}\times{\mathscr{C}}_{\ell}\times{\mathscr{C}}_{d}\times M^{2}$. Then, ${\mathcal{L}}^{\gamma}(X_{0},\alpha,\Xi^{\gamma,\epsilon}_{T})$ is the measure induced by the map, $\omega\in\Omega^{\gamma}\mapsto(X_{0}(\omega),\alpha_{1}(\omega),..,\alpha_{m}(\omega),\Xi^{\gamma,\epsilon}_{T}(\omega))\in{\mathscr{Y}}$. Since the state process $X^{\alpha,\epsilon}$ is a function of the differentials of $({\mathscr{W}}_{t},{\mathscr{N}}_{t})$ for $t\in[\epsilon,T]$, the independent increment property of these processes imply that law of $X^{\alpha,\epsilon}$ depends only on ${\mathcal{L}}^{\gamma}(X_{0},\alpha,\Xi^{\gamma,\epsilon}_{T})$. Then, the statement follows directly from the definition of $J_{g}^{\gamma,\epsilon}$. ∎
@@
Suppose Assumption 7.1 hold. Then, for any basis $\gamma$, $a_{0}\in U$, initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$ and control $\alpha\in{\mathscr{A}}(\gamma)$, there exists $\alpha^{k}\in{\mathscr{A}}_{s}(\gamma,a_{0})$ such that with $\epsilon_{k}:=t_{1}(\alpha^{k})$,
@@
In the absence of jumps, this result is proved in [26, Lemma 2], and we follow its proof closely. We first fix $\alpha\in{\mathscr{A}}(\gamma)$, and construct a sequence of simple controls $\alpha^{k}$ approximating $\alpha$ in the ${\mathscr{L}}^{2}(\mathrm{d}\mathbb{P}\otimes\mathrm{d}t)$-norm, [50, Lemma 4.4]. Then, the same estimates as in Lemma 7.6 show that $\lim_{k\to\infty}\mathbb{E}[\|X^{\alpha^{k},\epsilon^{k}}-X^{\alpha}\|_{T}^{2}]=0$. Following the proof of the aforementioned lemma, we use the continuity and boundedness of the functions $F,G$ to conclude that the cost functionals converge. ∎
@@ 7.4 Strong Controls
We next construct a strong control for each simple control in $\gamma$, by utilizing a classical stochastic transfer lemma, see for instance [40, Theorem 6.10].
@@
Suppose that $(S,{\mathcal{S}})$ is a measurable space and $(T,{\mathcal{T}})$ is a Borel space. Let $Y$ and $Z$ be random variables on a probability space $(\Omega,{\mathscr{F}},\mathbb{P})$ taking values in $S$ and $T$, respectively. Then, for any uniform random variable $U\in[0,1]$ independent of $X$, there exists a measurable function $\psi:[0,1]\times S\to T$ such that ${\mathcal{L}}^{\mathbb{P}}(Y,\psi(U,Y))={\mathcal{L}}^{\mathbb{P}}(Y,Z)$.
The following is the main technical step of the proof, and it is inspired by [26, Lemma 3].
@@
$t_{1}(\alpha^{st})=\epsilon,\qquad{\text{and}}\qquad J^{\gamma,\epsilon}(X_{0},\alpha)=J^{\gamma,\epsilon}(X_{0},\alpha^{st}).$
@@
We fix the basis $\gamma$, $X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)$, $a_{0}\in U$, and $\alpha\in{\mathscr{A}}_{s}(\gamma,a_{0})$. With $\alpha$ as in (7.4), set $\epsilon:=t_{1}(\alpha)$ and
@@
It is classical that there is a measurable map $\varphi:\mathbb{R}^{d}\mapsto[0,1]^{m}$ such that components of $\varphi(W^{\gamma}_{\epsilon})=:\xi=:(\xi_{1},..,\xi_{m})$ are independent uniform variables on $[0,1]$. Additionally, $(X_{0},\Xi_{1},..,\Xi_{m})$ and $\xi$ are independent.
Step 1. By the stochastic transfer lemma with $Y=X_{0}$, $Z=\alpha_{1}$, and $U=\xi_{1}$, there exists a measurable $\psi_{1}:[0,1]\times\mathbb{R}^{n}\mapsto U$ such that ${\mathcal{L}}(X_{0},\psi(\xi_{1},X_{0}))={\mathcal{L}}(X_{0},\alpha_{1})$. Then, for $i=2,..,m$, we apply the transfer lemma with $Y=(X_{0},\alpha_{1},..,\alpha_{i-1},\Xi_{i})$, $Z=\alpha_{i}$, and $U=\xi_{i}$ to construct a measurable map $\psi_{i}:[0,1]\times U^{i-1}\times\mathbb{R}^{n}\times{\mathscr{X}}\mapsto U$ such that
@@
Then, by Lemma 7.8, $J^{\gamma,\epsilon}(X_{0},\alpha)=J^{\gamma,\epsilon}(X_{0},\alpha^{st})$, and $\|\alpha\|_{2,T}=\|\alpha^{st}\|_{2,T}$. ∎
@@ Appendix A Poisson Point Processes
A function $p:D(p)\subset(0,\infty)\to E$ is called a point function on $E$, if its domain $D(p)$ is a countable and $\{s\in D(p)\cap(0,t]:p(s)\in S\}$ finite for every $t>0$ and a compact $S\subset E$. For every point function on $E$ we associate a counting measure $N_{p}$ on $[0,\infty)\times E$ by,
$N_{p}(\mathrm{d}t,\mathrm{d}\zeta)=\sum_{s\in D_{p}}\delta_{\{(s,p(s))\}}(\mathrm{d}t,\mathrm{d}\zeta).$
@@ Appendix B Discounted Infinite Horizon
Suppose that Assumptions 2.3, 7.1 hold. Then, for any $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$,

### v2

@@ 7.2 Delayed Start
This section introduces a technique presented in [28]. We now fix a stochastic basis $\gamma$, an initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$, and again suppress the dependence on $\gamma$. For $\alpha\in{\mathscr{A}}$ and $\epsilon>0$, we construct an approximating state process $X^{\alpha,\epsilon}$ and its cost. We first set
@@
$X^{\alpha,\epsilon}_{t}=X_{0}+\int_{\epsilon}^{t}$ $b(s,X^{\alpha,\epsilon}_{\cdot\wedge s},\mu^{\alpha,\epsilon}_{s},\alpha_{s})\,\mathrm{d}s+\int_{\epsilon}^{t}\boldsymbol{\sigma}(s,X^{\alpha,\epsilon}_{\cdot\wedge s},\mu^{\alpha,\epsilon}_{s},\alpha_{s})\,\mathrm{d}{\mathscr{W}}_{s}$
@@
where $\mu^{\alpha,\epsilon}_{s}={\mathcal{L}}(X^{\alpha,\epsilon}_{\cdot\wedge s}\,|\,{\mathscr{G}}_{s})$. Using ${\mathcal{L}}^{\alpha,\epsilon}_{t}:={\mathcal{L}}((X^{\alpha,\epsilon}_{\cdot\wedge t},\alpha_{t})\,|\,{\mathscr{G}})$, we define the approximating cost functional by,
@@
We first fix $\alpha\in{\mathscr{A}}$, $\epsilon>0$, and we suppress the dependence on $\mathbb{E}|X_{0}|^{2}$, and on the constants in Assumptions 2.3 and 7.2. To simplify the notation, we set
@@
$\mathbb{E}[\|X^{\alpha}-X^{\alpha,\epsilon}\|^{2}_{\epsilon}]=\mathbb{E}[\sup_{t\in[0,\epsilon]}|X^{\alpha}_{t}-X_{0}|^{2}]\leq c_{1}(\epsilon+\|\alpha\|^{2}_{2,\epsilon}).$
We then use the Burkholder-Davis-Gundy inequality, [45, Corollary 2.12], the Lipschitz property of the coefficients $(b,\boldsymbol{\sigma},\boldsymbol{\lambda})$, and the Grönwall’s inequality to arrive at
@@
$|F(t,X^{k}_{\cdot\land t},\alpha^{k}_{t},{\mathcal{L}}^{k}_{t})-F(t,Y^{k}_{\cdot\land t},\alpha^{k}_{t},\widehat{\mathcal{L}}^{k}_{t})|+|G(X^{k}_{\cdot\land T},{\mathcal{L}}^{k}_{T})-G(Y^{k}_{\cdot\land T},\widehat{\mathcal{L}}^{k}_{T})|\leq\boldsymbol{m}(\zeta_{k}),$
for every $t\in[0,T]$, where $\boldsymbol{m}$ is the modulus of continuity of $F$ and $G$. Consequently, the following sequence of random variables
@@ 7.3 Simple Controls
for some $t_{0}=0<t_{1}=:t_{1}(\alpha)<..<t_{m+1}=T$, ${\mathscr{F}}_{t_{i}}^{\gamma}$-measurable $\alpha_{i}:\Omega^{\gamma}\to U$ and $a_{0}\in U$.
@@
${\mathcal{L}}^{\gamma}(X_{0},\alpha,\Xi^{\gamma,\epsilon}_{T})={\mathcal{L}}^{\tilde{\gamma}}(\tilde{X}_{0},\tilde{\alpha},\Xi^{\tilde{\gamma},\epsilon}_{T})\quad\Rightarrow\quad J^{\gamma,\epsilon}(X_{0},\alpha)=J^{\tilde{\gamma},\epsilon}(\tilde{X}_{0},\tilde{\alpha}).$
@@
Consider the measurable space ${\mathscr{Y}}:=\mathbb{R}^{n}\times U^{m}\times{\mathscr{C}}_{\ell}\times{\mathscr{C}}_{d}\times M^{2}$. Then, ${\mathcal{L}}^{\gamma}(X_{0},\alpha,\Xi^{\gamma,\epsilon}_{T})$ is the measure induced by the map, $\omega\in\Omega^{\gamma}\mapsto(X_{0}(\omega),\alpha_{1}(\omega),..,\alpha_{m}(\omega),\Xi^{\gamma,\epsilon}_{T}(\omega))\in{\mathscr{Y}}$. Since the state process $X^{\alpha,\epsilon}$ is a function of the differentials of $({\mathscr{W}}_{t},{\mathscr{N}}_{t})$ for $t\in[\epsilon,T]$, the independent increment property of these processes implies that the law of $X^{\alpha,\epsilon}$ depends only on ${\mathcal{L}}^{\gamma}(X_{0},\alpha,\Xi^{\gamma,\epsilon}_{T})$. Then, the statement follows directly from the definition of $J^{\gamma,\epsilon}$. ∎
@@
Suppose Assumption 7.2 holds. Then, for any basis $\gamma$, $a_{0}\in U$, initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$ and control $\alpha\in{\mathscr{A}}(\gamma)$, there exists $\alpha^{k}\in{\mathscr{A}}_{s}(\gamma,a_{0})$ such that with $\epsilon_{k}:=t_{1}(\alpha^{k})$,
@@
In the absence of jumps, this result is proved in [29, Lemma 2], and we follow its proof closely. We first fix $\alpha\in{\mathscr{A}}(\gamma)$, and construct a sequence of simple controls $\alpha^{k}$ approximating $\alpha$ in the ${\mathscr{L}}^{2}(\mathrm{d}\mathbb{P}\otimes\mathrm{d}t)$-norm, [53, Lemma 4.4]. Then, the same estimates as in Lemma 7.7 show that $\lim_{k\to\infty}\mathbb{E}[\|X^{\alpha^{k},\epsilon^{k}}-X^{\alpha}\|_{T}^{2}]=0$. Following the proof of the aforementioned lemma, we use the continuity and boundedness of the functions $F,G$ to conclude that the cost functionals converge. ∎
@@ 7.4 Strong Controls
We next construct a strong control for each simple control on $\gamma$, by utilizing a classical stochastic transfer lemma, see for instance [44, Theorem 6.10].
@@
Suppose that $(S,{\mathcal{S}})$ is a measurable space and $(T,{\mathcal{T}})$ is a Borel space. Let $Y$ and $Z$ be random variables on a probability space $(\Omega,{\mathscr{F}},\mathbb{P})$ taking values in $S$ and $T$, respectively. Then, for any uniform random variable $V\in[0,1]$ independent of $Y$, there exists a measurable function $\psi:[0,1]\times S\to T$ such that ${\mathcal{L}}^{\mathbb{P}}(Y,\psi(V,Y))={\mathcal{L}}^{\mathbb{P}}(Y,Z)$.
The following is the main technical step of the proof, and it is inspired by [29, Lemma 3].
@@
$t_{1}(\alpha^{st})=\epsilon\quad{\text{and}}\quad J^{\gamma,\epsilon}(X_{0},\alpha)=J^{\gamma,\epsilon}(X_{0},\alpha^{st}).$
@@
We fix $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$, $a_{0}\in U$, and $\alpha\in{\mathscr{A}}_{s}(\gamma,a_{0})$. With $\alpha$ as in (7.5), set $\epsilon:=t_{1}(\alpha)$ and
@@
It is classical that there is a measurable map $\varphi:\mathbb{R}^{d}\to[0,1]^{m}$ such that components of $\varphi(W^{\gamma}_{\epsilon})=:\xi=:(\xi_{1},..,\xi_{m})$ are independent uniform variables on $[0,1]$. Additionally, $(X_{0},\Xi_{1},..,\Xi_{m})$ and $\xi$ are independent.
Step 1. By the stochastic transfer lemma with $Y=X_{0}$, $Z=\alpha_{1}$, and $V=\xi_{1}$, there exists a measurable $\psi_{1}:[0,1]\times\mathbb{R}^{n}\to U$ such that ${\mathcal{L}}(X_{0},\psi_{1}(\xi_{1},X_{0}))={\mathcal{L}}(X_{0},\alpha_{1})$. Then, for $i=2,..,m$, we apply the transfer lemma with $Y=(X_{0},\alpha_{1},..,\alpha_{i-1},\Xi_{i})$, $Z=\alpha_{i}$, and $V=\xi_{i}$ to construct a measurable map $\psi_{i}:[0,1]\times\mathbb{R}^{n}\times U^{i-1}\times{\mathscr{X}}\to U$ such that
@@
Then, by Lemma 7.9, $J^{\gamma,\epsilon}(X_{0},\alpha)=J^{\gamma,\epsilon}(X_{0},\alpha^{st})$ and $\|\alpha\|_{2,T}=\|\alpha^{st}\|_{2,T}$. ∎
@@ Appendix A Poisson Point Processes
A function $p:D(p)\subset(0,\infty)\to E$ is called a point function on $E$, if its domain $D(p)$ is countable and $\{s\in D(p)\cap(0,t]:p(s)\in S\}$ finite for every $t>0$ and a compact $S\subset E$. For every point function on $E$ we associate a counting measure $N_{p}$ on $[0,\infty)\times E$ by,
$N_{p}(\mathrm{d}t,\mathrm{d}\zeta)=\sum_{s\in D(p)}\delta_{\{(s,p(s))\}}(\mathrm{d}t,\mathrm{d}\zeta).$
@@ Appendix B Discounted Infinite Horizon
Suppose that Assumptions 2.3, 7.2 hold. Then, for any $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$,